# PPFL-RDSN: Privacy-Preserving Federated Learning-based Residual Dense Spatial Networks for Encrypted Lossy Image Reconstruction

### v1

@@ 6 Conclusion
In this paper, we have presented an innovative Privacy-Preserving Federated Learning (PPFL) framework for training Residual Dense Spatial Networks (RDSN) and designed a multi-party Encryption-the-Lossy-Compression scheme based on [21]. Our experimental results show the PPFL-RDSN framework’s advantages in maintaining high-quality and robust lossy image reconstruction while decreasing computational overhead and enhancing privacy and security. The clients perform their local differentially private training to obtain the initial model. After aggregation and model integrity verification, they execute the multi-party EtLC scheme and determine whether the initial model demands refinement iteratively. We show that the PPFL-RDSN training system is scalable to a large client pool and provides strong privacy guarantees offered by LDP and protection against model poisoning or substitution by untrusted aggregators. Our framework makes an important contribution in employing privacy-preserving federated learning for image processing tasks, paving the way for more privacy-protecting, secure, and efficient collaborative learning solutions essential for data-intensive applications with distributed datasets. Future research directions can focus on optimizing the RDSN structures to improve the model performance further, incorporating other privacy protection mechanisms to address a broader set of privacy and security threats to enhance the framework. Also, exploring hybrid approaches that combine multiple privacy techniques, such as DP and multiparty computation, is a potentially promising research direction to address practical challenges for real-world applications.

### v3

@@ V Conclusion and Discussion
In this paper, we have presented an innovative Privacy-Preserving Federated Learning (PPFL) framework for training Residual Dense Spatial Networks (RDSN) and designed a multi-party Encryption-then-Lossy-Compression scheme based on [21]. Our experimental results show the proposed PPFL-RDSN framework’s advantages in maintaining high-quality and robust lossy image reconstruction while enhancing privacy and security. The clients perform their local differentially private training to obtain the initial model. After aggregation and model integrity verification, they execute the multi-party EtLC scheme and determine whether the initial model demands refinement iteratively. We show that the PPFL-RDSN training system is scalable to a larger client pool and provides strong privacy guarantees offered by LDP and protection against model poisoning or substitution by untrusted aggregators. Our framework makes an important contribution in employing privacy-preserving federated learning for image processing tasks, paving the way for more privacy-protecting, secure, and efficient collaborative learning solutions essential for data-intensive applications with distributed datasets. Future research directions can focus on optimizing the RDSN structures to improve the model performance further, incorporating other privacy protection mechanisms to address a broader set of privacy and security threats to enhance the framework. Also, exploring hybrid approaches that combine multiple privacy techniques, such as DP and multiparty computation, is a potentially promising research direction to address practical challenges for real-world applications.
Limitations. Our design emphasizes a practical baseline—high-frequency DCT perturbation with fixed $(\tau,C_{\max})$ and moments-accountant composition, a sparse-index watermark with standard key management, and an honest-but-curious multi-party EtLC prototype with FedAvg; adaptive frequency protection, richer collusion models, and secure or Byzantine-robust aggregation are natural extensions. Empirically, we evaluate on canonical SR datasets with a 10-client simulation and representative $(\epsilon,\delta)$; scaling to larger and heterogeneous client pools and wide-area settings, broadening to non-vision tasks, and expanding robustness and $\epsilon$–utility sweeps are promising directions.